OpenStudy (anonymous):

‎1.what is the value of K if the line 3x-2y=1 is perpendicular to the line y-(2K+1)x=3? 2.write an equation of the line in S.F. a.having a slope of 2/3 and a y-intecept of -5 b.having a slope of -5 and passing through (2,-5) c.with an x-intercept of -5 and a y-intercept of -3 d.parallel to y=2x-3 and passing through (3,-1) e.perpendicular to y+1=3x and passing through (2,-3) 3. find y if the line through A(-1,2) and B(3,5) is parallel to the line through (0,-1) and (6,y)

OpenStudy (anonymous):

when two lines are perpendicular to each other their slopes are related as m1 = -1/m2 now you have two equations 3x-2y=1 and y-(2K+1)x=3 writing these equations in term of x we get can you do it???

OpenStudy (anonymous):

I can't Please teach me !

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

take the first equation

OpenStudy (anonymous):

3x-2y=1 subtracting 3x from both sides we get 3x-3x-2y = 1-3x -2y = 1-3x now dividing both sides by -2 we get y = 3/2x-1/2 got it???

OpenStudy (anonymous):

Yes ! Then ? (:

OpenStudy (anonymous):

now consider the second equation

OpenStudy (anonymous):

y-(2K+1)x=3 adding (2K+1)x to both sides we get y = (2K+1)x +3 clear???

OpenStudy (anonymous):

then ?

OpenStudy (anonymous):

now you have these both equations y = 3/2x-1/2 and y = (2K+1)x +3 these both equations are in slope intercept form as we know that the general form of slope intercept form is y = mx+b here m = slope and b = y-intercept

OpenStudy (anonymous):

can you guess the slopes of these both lines now??

OpenStudy (anonymous):

wait . How about the K ?

OpenStudy (anonymous):

we are getting to it but first find the slope

OpenStudy (anonymous):

y = 3/2x-1/2 what is the slope now??

OpenStudy (anonymous):

3/2

OpenStudy (anonymous):

yes exactly

OpenStudy (anonymous):

then ?

OpenStudy (anonymous):

y = (2K+1)x +3 and the here the slope is ??

OpenStudy (anonymous):

guess it we are one step away from finding K

OpenStudy (anonymous):

2k+1 ?

OpenStudy (anonymous):

hmm thats right

OpenStudy (anonymous):

now these line are perpendicular to each other thus their slopes are related as m1 = -1/m2 here m1 = 3/2 and m2 = 2k+1 3/2 = -1/(2k+1) can you solve for K now???

OpenStudy (anonymous):

K=-1 ?

OpenStudy (anonymous):

how you get that??

OpenStudy (anonymous):

3/2=-1/2k+1 6k+3=-2 6k=-2-3 6k=-6 divided by 6 k=-1

OpenStudy (anonymous):

no no its -5 -2-3 = -5

OpenStudy (anonymous):

aw . hah ! Sorry XD

OpenStudy (anonymous):

so K = -5/6 is the required value of K

OpenStudy (anonymous):

i think so ^^

OpenStudy (anonymous):

it is the right answer verify it from the book

OpenStudy (anonymous):

Yeah ! How about in number 2 ?

OpenStudy (anonymous):

now you have y-(2K+1)x=3 put in the value of K you got

OpenStudy (anonymous):

do the last step and you are done

OpenStudy (anonymous):

Done ! Correct ^^

OpenStudy (anonymous):

are you sure??

OpenStudy (anonymous):

how about letter d in number 2 ?

OpenStudy (anonymous):

Yes ! 100% sure ^_^

OpenStudy (anonymous):

d.parallel to y=2x-3 and passing through (3,-1) as two lines are parallel their slopes are equal thus the slope from the given equation y=2x-3 is ???

OpenStudy (anonymous):

find it

OpenStudy (anonymous):

the slope is 2

OpenStudy (anonymous):

ah ! I Get IT ^^

OpenStudy (anonymous):

yes so the equation of line you want to find will be having the same slope of 2 now you have slope and one point (3,-1) use point slope form to get the equation

OpenStudy (anonymous):

yes ! haha .. I get it ^^ New Achievement XD

OpenStudy (anonymous):

but one mistake which i did is part a

OpenStudy (anonymous):

In letter E ?

OpenStudy (anonymous):

basically this was the question and by mistake i told you to put value of K in the equation y-(2K+1)x=3 so that was incorrect, so be sure you have done this part correctly 2.write an equation of the line in S.F. a.having a slope of 2/3 and a y-intecept of -5

OpenStudy (anonymous):

y=2/3x-5 isolate the 3 . Multiply both sides 3y=2x-15 -2x+3y=-15 multiply by -1 2x-3y=15

OpenStudy (anonymous):

yes yes you are right

OpenStudy (anonymous):

now go for part e

OpenStudy (anonymous):

e.perpendicular to y+1=3x and passing through (2,-3) as i told you when two lines are perpendicular their slopes are related as m1 = -1/m2 so what is the slope from this equation y+1=3x

OpenStudy (anonymous):

the slope is 3

OpenStudy (anonymous):

yes now the slope of the line will will be perpendicular to it will be ???

OpenStudy (anonymous):

Gets !

OpenStudy (anonymous):

*which

OpenStudy (anonymous):

as you have m1 = 3 m1 = -1/m2 now find m2

OpenStudy (anonymous):

m2 is -1/3 the equation is x+3y=-4

OpenStudy (anonymous):

use point slope form to get the equation

OpenStudy (anonymous):

yap :)))

OpenStudy (anonymous):

Yes ! Thank you so much ^^

OpenStudy (anonymous):

i m sure your answer is correct

OpenStudy (anonymous):

you are welcome and forgive my mistakes

OpenStudy (anonymous):

Thanks !!! Haha ...

OpenStudy (anonymous):

you are welcome

OpenStudy (anonymous):

:))

OpenStudy (anonymous):

Oh ! In number 3 :)

OpenStudy (anonymous):

find y if the line through A(-1,2) and B(3,5) is parallel to the line through (0,-1) and (6,y)

OpenStudy (anonymous):

then ?

OpenStudy (anonymous):

first consider these points can you find the slope from these points?? A(-1,2) and B(3,5)

OpenStudy (anonymous):

\[slope = \frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (anonymous):

then ?

OpenStudy (anonymous):

so what is the slope???

OpenStudy (anonymous):

3/4

OpenStudy (anonymous):

yes now consider the next two points (0,-1) and (6,y) can you find the slope??

OpenStudy (anonymous):

same slope because it's parallel :))

OpenStudy (anonymous):

yes but here what is the slope you got??

OpenStudy (anonymous):

3/4

OpenStudy (anonymous):

no no from the nest two points containing y

OpenStudy (anonymous):

*next

OpenStudy (anonymous):

the value of y is 7/2 ^^

OpenStudy (anonymous):

(0,-1) and (6,y) slope = (y+1)/6 now equate these slopes (y+1)/6 = 3/4 y+1 = 18/4 y = 18/4 - 1

OpenStudy (anonymous):

I used the equation slope=y2-y1/x2-x1 3/4=y+1/6-0 3/4=y+1/6 cross multiply 4y+4=18 4y=18-4 divided by 4 y=14/4 Simplest Form y=7/2

OpenStudy (anonymous):

yes yes thats right

OpenStudy (anonymous):

by the way it seems to be like you are my teacher

OpenStudy (anonymous):

haha ^^

OpenStudy (anonymous):

keep your good and fast and accurate mathematics up

OpenStudy (anonymous):

Thanks ^^

OpenStudy (anonymous):

Thanks for the Time ^______^

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

:))